ملاحظة: اجب على سؤال واحد فقط Q5: prove that the L{ear} 3 = 6-a) by definition of 1 f Laplace transform ?
Answer #1

Ans. Prove (fear} auta Laplace transform definition says that L{f(t) -st e f(t) at F (S) - ax Now Heue the function is givenjest, eat at Now this can be written al ->) - s eat-st dt 2 Now let U= at-st. differentiate on both side with respect to t soat -st S-a Now apply the limits so we get 8 atast = ) e [ Q-g =) alco)-S (0) e a(o)-sco) e a-S a - S a - S Note : a (or) – sl*[-(+) [: =) a-S = ) ㅗ S-a f(s) Ans. Hence Proved.